2. (40 marks) The following equations have been estimated by OLS using annual data for a Canadian city for 1991-2018. QW per capita consumption of wine in litres RPW = relative price of wine RPB relative price of beer RPS relative price of spirits RY real per capita disposable income (1) QW- 10.50-7.30RPW + 0.003RY 4.96/(2.00) 0.0002) R2= 0.90 a2= 3.50 2) QW.00H6.9ORPW 0.0050RY 1.10RPB 1.50RPS 10.40 2.30 (0.0009) (7.65) (0.55) (a) (10 marks) Perform a test of the significance of the regression on equation (1) using 1 % level of significance. (b) (10 marks) Using equation (2) determine whether the cross price elasticity of demand for wine is elastic or inelastic with respect to the price of beer. Briefly explain the meaning of this cross elasticity. (Note that QW-6.00 and RPB-0.55) (c) (10 marks) Perform a test on the null hypothesis that the coefficients on RPB and RPS are jointly equal to zero using 5% level of significance (d) (10 marks) In 1997 RY-4200, RPW-1.50, RPB-0.5, RPS 5.3. As part of its 2019 budget the local government proposes to impose a 5% consurnption tax on wine. Assuming that the beer and spirits producers are not planning to change prices and that real income is expected to grow by 5% what is your forecast of the wine consumption in 2019 (Use equation 2). Please answer all four subquestions with specific steps
Answer #1

Consider the following notations: QW per capita consumption of wine in liters RPW relative price of wine RPB relative price oHere, as follows MSR is the regression mean square and is given SSR MSR k-1 And MSE is the error mean square which is given aThe value of the test statistic is computed as follows 28-3 F = 3 -11-0.90 0.90 12.5 (9) =112.5 Now the tabulated value of F(c) as follows The hypothesis for testing that the coefficients on RPB and RPS is given Null Hypothesis, Ho B B4 =0 AlternatiFor equation (2), 6 =2.50 and c = MSE Therefore (28 -5)V2.5 = toresti cted SSE 36.3662 1oresticted Thus, the test statistic i= 4410 The forecast for wine consumption using equation (2) is obtained as follows: QW -4.00 6.90(1.575)+0.005 (4410)+1.1(0.5